$SPAD/src/input bug6357.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool bug6357.output
)set message test on
)set message auto off
)clear all
-- The original author assumed (roughly) that sqrt(1/x)=1/sqrt(x),
-- which is wrong (for example,
-- sqrt(-1/2) = %i/sqrt(2) != 1/(%i*sqrt(2)) = -%i/sqrt(2)
--S 1 of 2
sqrt(-1/2)
--R
--R
--R         +---+
--R        \|- 1
--R   (1)  ------
--R          +-+
--R         \|2
--R                                                        Type: AlgebraicNumber
--E 1
--S 2 of 2
sqrt(-1/abs(x))-1/sqrt(-abs(x))
--R
--R
--R                    +--------+
--R         +--------+ |     1
--R        \|- abs(x)  |- ------  - 1
--R                   \|  abs(x)
--R   (2)  --------------------------
--R                 +--------+
--R                \|- abs(x)
--R                                                    Type: Expression(Integer)
--E 2
)spool
)lisp (bye)

———-

References^{[1] nothing}

4